Goodwin Technologies, a relatively young company, has been wildly successful but has yet to pay a dividend. An analyst forecasts that Goodwin is likely to pay its first dividend three years from now. She expects Goodwin to pay a $1.50000 dividend at that time (D₃ = $1.50000) and believes that the dividend will grow by 7.80000% for the following two years (D₄ and D₅). However, after the fifth year, she expects Goodwin’s dividend to grow at a constant rate of 3.42000% per year.

Goodwin’s required return is 11.40000%. Fill in the following chart to determine Goodwin’s horizon value at the horizon date (when constant growth begins) and the current intrinsic value. To increase the accuracy of your calculations, do not round your intermediate calculations, but round all final answers to two decimal places.

Term

Value

Horizon value   
Current intrinsic value   

Assuming that the markets are in equilibrium, Goodwin’s current expected dividend yield is   , and Goodwin’s capital gains yield is   .

Goodwin has been very successful, but it hasn’t paid a dividend yet. It circulates a report to its key investors containing the following statement:

Goodwin has yet to record a profit (positive net income).

Is this statement a possible explanation for why the firm hasn’t paid a dividend yet?

Yes

No
Answer #1

Rate positively ..

ans a Computation of horizon value at the end of 5 year = Dividend for 6th year / (required rate - growth rate)
year dividend
3 $     1.50
4 $     1.62 1.5*107.8%
5 $     1.74 1.62*107.8%
6 $     1.80 1.74*103.42%
therefore horizon value = =1.8/(11.4%-3.42%)
      22.59
ans b Current intrensic value = Present value of future cash flow @ 11.4% rate of return
i ii iii=i+ii iv v=iii*iv
year dividend horizon value Total cash flow PVIF @ 11.4% present value
3 $     1.50          1.50 0.7233432 $                              1.09
4 $     1.62          1.62 0.6493207 $                              1.05
5 $     1.74       22.59       24.33 0.5828731 $                            14.18
$                            16.32
therefore intrensic value= $                            16.32
ans c Expected dividend at the end of year/Price at the beginning of year
Price in beginning of year - 1
i ii iii=i+ii iv v
year dividend horizon value Total cash flow PVIF @ 11.4% present value
3 $     1.50          1.50 0.8058044 $                              1.21
4 $     1.62          1.62 0.7233432 $                              1.17
5 $     1.74       22.59       24.33 0.6493207 $                            15.80
$                            18.18
Dividend yield = 0/18.18 0.00%
Capital gain yield = 11.4%-0 12.84%
ans d Correct answer is option : YES